# Distributed Stochastic Gradient Descent with Staleness: A Stochastic Delay Differential Equation Based Framework

### v1

@@ III-B The Convergence Speed Given Gradient Noise
In this subsection, we derive a relationship between the convergence time and the gradient noise. To gain more insight on the impact of the gradient noise on the convergence time, we investigate the stochastic processes $\mathrm{d}x_{i}(t)=-v_{i}x_{i}(t)\mathrm{d}t+\sigma\mathrm{d}B(t)$ for $i=1,2,...,N$ in this subsection. In this way, the convergence time of SGD can modeled as a first hitting time of (5) to the $\delta$-neighborhood of the global optimum $\boldsymbol{x}^{*}$, i.e.,
$H=\inf\{t>0:\left\|{\boldsymbol{x}}(t)-\boldsymbol{x}^{*}\right\|_{\infty}<\delta\},$ (9)
in which the $i$th element of $\boldsymbol{x}^{*}$ is denoted by $x_{i}^{*}$. Defining $H_{i}=\inf\{t>0:|x_{i}(t)-x_{i}^{*}|<\delta\}$ for $i=1,2,...,N$, it holds that $H=\max\limits_{i}\{T_{i}\}$. [^6] In this paper, it is generally assumed that the SGD algorithm doesn’t leave the $\delta$-neighborhood of the global optimum once it enters it when the SGD algorithm converges. To further demonstrate the impact of the gradient noise on the convergence time of SGD, in this subsection, we first introduce the following lemma on the first hitting time according to [50]. [^7] For example, the fact that the expected hitting time of a given level by a standard Brownian motion is infinite is a direct corollary of Lemma 1.
Lemma 1
The expectation of the first hitting time $H_{L}=\inf\{t>0:x_{1}\leq x(t)\leq x_{1}+\mathrm{d}x,x(0)=x_{0}\}$ of the one-dimensional stochastic process $\mathrm{d}x(t)=s(x)\mathrm{d}t+\sigma(x)\mathrm{d}B(t)$ is $\mathbb{E}\{H_{L}\}=\int_{x_{0}}^{x_{1}}\frac{1}{s(x)}\mathrm{d}x,$ in which $s(x)$ is the drift velocity.
By Lemma 1, it holds that
@@
The probability distributions of the first hitting times $H_{i}$ are however complex. To shed more light on the probability characteristics of $H$, we shall next present an approximation of the probability distribution of $T_{i}$. Since the first moment of $H_{i}$ has already been revealed by Eq. (10), we are interested in the second moment of $T_{i}$ in the following discussion.
Given an OU process described by Eq. (6), $x_{i}(t)$ follows a Gaussian distribution with mean $\mathbb{E}\{x_{i}(t)\}=x_{i,0}e^{-v_{i}t}$ and variance $\text{Var}\{x_{i}(t)\}=\frac{\sigma^{2}}{2v_{i}}\left(1-e^{-2v_{i}t}\right)$. The drift velocity at the time that the OU process enters the $\delta$ neighborhood of the global optimum can be approximated by $s\approx v_{i}\delta$ by Eq. (6). By this means, the variance of the first hitting time $H_{i}$ can be given by $\text{Var}\{H_{i}\}\approx\frac{\sigma^{2}}{2v_{i}^{3}\delta^{2}}\left(1-\frac{1}{\alpha_{i}^{2}}\right)$, in which $\alpha_{i}=x_{i,0}/\delta$. Then, the first hitting time can be approximated as Gaussian distributed as follows:
@@
Next, we provide further discussion on the probability distribution of $H=\max_{i}\{H_{i}\}$. Since the first hitting times are approximated Gaussian distributed, we have $\mathbb{E}\{H\}\approx E\{H_{i}\}$ if $E\{H_{i}\}\gg E\{H_{k}\}$ for $k\neq i$ in which $i=\arg\max_{j}E\{H_{j}\}$.
Further, $H_{i}$ are identically distributed random variable when $\boldsymbol{V}=v\boldsymbol{I}$ and $\hat{\boldsymbol{x}}(0)=c\boldsymbol{1}$, in which $v>0$, $\boldsymbol{I}$ denotes the identity matrix, and $\boldsymbol{1}$ denotes the vector of ones. In this case, we have $\mathbb{E}\{H_{i}\}=\frac{1}{v}\log\alpha$ in which $\alpha=\frac{c}{\delta}$ and $\text{Var}\{H_{i}\}\approx\frac{\sigma^{2}}{2v^{3}\delta^{2}}\left(1-\frac{1}{\alpha^{2}}\right)$. Let
@@
in which $\Xi_{N}=\max\{Z_{1},Z_{2},...,Z_{N}\}$ and $Z_{i}$s are independent and identically distributed (i.i.d.) standard Gaussian distributed random variable $Z_{i}\sim\mathcal{N}(0,1)$. More specifically, $\Xi_{N}$ is the maximum of $N$ i.i.d. standard Gaussian random variables. In addition, $\Theta(N)$ is on the order of $\mathcal{O}(1)$. In particular, $\Theta(1)=0$. In this way, it holds that
$\mathbb{E}\{H\}=\frac{1}{v}\log\alpha+\frac{\sigma\sqrt{1-\frac{1}{\alpha^{2}}}}{\sqrt{2v^{3}}\delta}\left(\sqrt{\log N}+\Theta(N)\right),$ (13)
@@
When $N=1$, it has been shown that the gradient noise does not affect the expectation of the first hitting time in the SGD algorithm by Lemma 1. While Eq. (13) also suggests the same conclusion. Further, the presence of the gradient noise induced an additional term on the average first hitting time, which is propositional to the standard variance of the gradient noise and $\sqrt{\log N}$ when $N\geq 2$.
@@
Next, in order to make the problem analytically tractable and get insightful results, we consider the impact of gradient noise with $f(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{I}\boldsymbol{x}$. Thus, it holds that $\breve{\boldsymbol{x}}[j+1]=\breve{\boldsymbol{x}}[j]-\eta\breve{\boldsymbol{x}}[j]+\boldsymbol{w},$ in which $\eta$ is the learning rate and $\boldsymbol{w}$ is an isotropic Gaussian noise in which $\boldsymbol{w}\sim\mathcal{N}(0,\breve{\sigma}^{2}\boldsymbol{I})$. In this case, we define $\boldsymbol{r}[j]=\breve{\boldsymbol{x}}[j]-\boldsymbol{x}^{*}$ and $r=\left\|\boldsymbol{r}\right\|_{2}$ to simplify the notations. Due to the symmetric of the normal distribution, the projection of $\boldsymbol{w}$ onto the direction of $\boldsymbol{r}$, denoted by $\boldsymbol{w}_{\parallel}$, is also Gaussian distributed, i.e., $\left\|\boldsymbol{w}_{\parallel}\right\|_{2}\sim\mathcal{N}(0,\breve{\sigma}^{2})$. In this context, the component perpendicular to the $\boldsymbol{r}$ direction is given by $\boldsymbol{w}_{\perp}=\boldsymbol{w}-\boldsymbol{w}_{\parallel}$. In this case, we have $\langle\boldsymbol{w}_{\perp},\boldsymbol{r}\rangle=0$, in which $\langle\boldsymbol{a},\boldsymbol{b}\rangle$ denotes the inner product of the two vector $\boldsymbol{a}$ and $\boldsymbol{b}$. Also, it can be seen that $\mathbb{E}\{\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}\}=(N-1)\breve{\sigma}^{2}$. Thus, it holds that
$r[j+1]=r[j]-\eta r[j]+\left(\sqrt{r^{2}[j]+\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}-r[j]\right)+w_{\parallel},$ (14)
@@
in which $w_{\parallel}=\left\|\boldsymbol{w}_{\parallel}\right\|_{2}$. For $r\gg\breve{\sigma}$, we have $\sqrt{r^{2}+\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}-r\approx\frac{\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}{2r}$, since the approximation $\sqrt{1+s}\approx 1+\frac{1}{2}s$ holds around the point $s=0$. For a given $r$, let the backward speed be defined as $z=\frac{\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}{2r}$ In this case, it can be seen that the presence of the noise $\boldsymbol{w}_{\perp}$ hinders the process of the gradient descent, the speed of which is given by $z$. More specifically, $z$ is Chi-square distributed with mean given by $\mathbb{E}\left\{z\right\}=\frac{(N-1)\breve{\sigma}^{2}}{2r}$, and variance given by $\text{Var}\left\{z\right\}=\frac{(N-1)\breve{\sigma}^{4}}{r^{2}}$. When $N=2$, the backward speed $z$ follows an exponential distribution. Therefore, the continuous-time approximation of Eq. (14) can be formulated as follows:
$\mathrm{d}r=-\left(\eta r-\frac{(N-1)\breve{\sigma}^{2}}{2r}\right)\mathrm{d}t+\sqrt{\sigma^{2}+\frac{(N-1)\breve{\sigma}^{4}}{r^{2}}}\mathrm{d}B(t).$ (15)
@@
The expectation of the first hitting time $H_{d}$ is given by $\mathbb{E}\{H_{d}\}=\mathcal{F}(r_{0})-\mathcal{F}(\delta)$, in which $F(r)=\frac{1}{2\eta}\log\left({\eta r^{2}-\frac{(N-1)\breve{\sigma}^{2}}{2}}\right)$.
@@
According to the primitive function of $f(r)={\left(\eta r-\frac{(N-1)\breve{\sigma}^{2}}{2r}\right)^{-1}}$ with respect to $r$ and Lemma 1, we arrive at the corollary. ∎
In this way, one can notice that the presence of the gradient noise does not affect the expectation of the first hitting time of the gradient descent algorithm when $N=1$. Further, the above analysis also indicates that as the value of $r$ decreases, the adverse impact of gradient noise on convergence becomes increasingly severe. This also suggests that as the training process progresses, increasing the number of users involved in training or enlarging the mini-batch size to reduce the power of gradient noise could accelerate the convergence of SGD. Note that a high-resolution quantization of the gradients with large gradient noise can lead to a lot of network bandwidth usage, thereby being inefficient due to the waste of communication resources.
@@ III-C The Convergence Speed Given Gradient Staleness
In this subsection, we present an analysis of the SGD with stale through an SDDE-based approach. For the sake of discussion, we shall first focus on the fixed staleness before extending our discussion to the case in which $\tau$ is a random variable. Let the gradient staleness in SDDE be denoted by a positive constant $\tau$. First, we focus on the characteristics of the solution to the deterministic differential equation. By substituting the sampling function $\varrho(t)=Ae^{{\lambda}t}$, in which ${\lambda}$ is a complex number into the delay differential equation (8), we have the following characteristics equation:
@@
Therefore, Eq. (16) has a single positive root when ${v}<0$. In this case, the SGD algorithm with a stale gradient overdamps and does not converge to the global minimum. Further, the characteristic equation has two negative real roots when $0<{v}\tau<\frac{1}{e}$. In this case, the SGD algorithm with a stale gradient decays monotonically to the global minimum of the quadratic objective function. Moreover, the characteristics equation has a pair of complex conjugate roots with negative real parts when $\frac{1}{e}<{v}\tau<\frac{\pi}{2}$. In this case, the SGD algorithm with stale gradient undergoes damped oscillations. When ${v}\tau>\frac{\pi}{2}$, the real parts of the roots of Eq. (16) becomes positive. Therefore, the SGD algorithm undergoes diverging oscillations. Note that $\eta$ is the step size of the SGD algorithm. According to the above discussion, we have the following theorem for the discrete-time SGD.
@@
The two roots of the characteristics function Eq. (16) are given by
@@
Given the learning rate $\eta$ and the step staleness $\breve{\tau}$, the characteristic roots of the continuous-time approximation of the asynchronous SGD algorithm, $\breve{x}[j+1]=\breve{x}[j]-\eta v\breve{x}\left[{j-\breve{\tau}}\right]$, are given by $\lambda_{k}=\frac{W_{k}{(-{v}\eta\breve{\tau})}}{\eta\breve{\tau}}$ for $k=0,1$.
@@
With a fixed step size, $\lambda_{k}$ are real and negative in which $|\lambda_{0}|\leq|\lambda_{1}|$ when $\breve{\tau}<\frac{1}{e\eta{v}}$. Besides, it can be noticed that $|\lambda_{0}|$ increases as $\breve{\tau}$ increases for $\breve{\tau}\in(0,\frac{1}{e\eta{v}})$. This indicates that the presence of a small degree of gradient staleness accelerates the stochastic gradient descent algorithm, which aligns with the analysis in [14]. However, the SGD algorithm starts to oscillate when $\breve{\tau}\in(\frac{1}{e\eta{v}},\frac{\pi}{2\eta{v}})$. In addition, the real part of the roots became positive when $\breve{\tau}>\frac{\pi}{2\eta{v}}$. It indicates that, the SGD algorithm diverges with a large degree of gradient staleness.
Remark:  To reveal the characteristic roots of (8) in case $\tau$ is a random variable, we will present the characteristic roots with uniformly distributed gradient staleness in the forthcoming discussion. Consider Gaussian distributed gradient staleness, i.e., $\tau\sim\mathcal{N}(\vartheta,\varsigma^{2})$. Empirically, we find that it holds that $\lambda_{k}\approx W_{k}(-v\vartheta)/\vartheta$ for $k=0,1$ when $\varsigma\ll\vartheta$ [^9] More specifically, we have the following observation of Gaussian distributed gradient staleness through numerical results. When $\tau\sim\mathcal{N}(\vartheta,\varsigma^{2})$ and $\varsigma\ll\vartheta$, asynchronous SGD nearly has the same convergence rate as when the gradient staleness is a constant, i.e., $\tau=\vartheta$. With a sufficiently large $\varsigma$, asynchronous SGD may diverge despite a small $\vartheta$. Analytical results for the staleness of Gaussian distribution gradients are noted as future work.

### v2

@@ III-B The Convergence Speed Given Gradient Noise
In this subsection, we derive a relationship between the convergence time and the gradient noise. To gain more insight on the impact of the gradient noise on the convergence time, we investigate the stochastic processes
$\mathrm{d}x_{i}(t)=-v_{i}x_{i}(t)\mathrm{d}t+\sigma\mathrm{d}B(t),$ (17)
for $i=1,2,...,N$ in this subsection. In this way, the convergence time of SGD can be modeled as a first hitting time of (11) to the $\delta$-neighborhood of the global optimum $\boldsymbol{x}^{*}$, i.e.,
$\tilde{H}=\inf\{t>0:\left\|{\boldsymbol{x}}(t)-\boldsymbol{x}^{*}\right\|_{\infty}<\delta\},$ (18)
in which the $i$th element of $\boldsymbol{x}^{*}$ is denoted by $x_{i}^{*}$. Defining
$H_{i}=\inf\{t>0:|x_{i}(t)-x_{i}^{*}|<\delta\},$ (19)
for $i=1,2,...,N$. Note that the definition of the first hitting time is presented in Subsection II-C. To further demonstrate the impact of the gradient noise on the convergence time of SGD, in this subsection, we first introduce the following proposition on the first hitting time according to [49].
Proposition 1
The expectation of the first hitting time of an OU process (17) is given by
@@
The probability distributions of the first hitting times $H_{i}$ are however complex. To shed more light on the probability characteristics of $H$, we shall next show that the first hitting time $T_{i}$ is approximately Gaussian distributed.
Proposition 2
The first hitting time (18) of the OU process (17) can be approximated by a Gaussian distributed random variable
@@
Proof:
See Appendix A. ∎
Next, let $H_{\max}=\max_{i}\{H_{i}\}$ be the maximum of the first hitting times. In this context, a careful reader may notice that the $H_{\max}$ is not necessarily identical to $H$, while the difference between $H_{\max}$ and $\tilde{H}$ will be detailed in Appendix B
Further, $H_{i}$s are identically distributed random variables when $\boldsymbol{V}=v\boldsymbol{I}$ and $\hat{\boldsymbol{x}}(0)=c\boldsymbol{1}$, in which $v>0$, $\boldsymbol{I}$ denotes the identity matrix, and $\boldsymbol{1}$ denotes the vector of ones. In this case, ${H_{\max}}$ is given by the expectation of maximum of i.i.d. Gaussian variables. Let
@@
in which $\Xi_{N}=\max\{Z_{1},Z_{2},...,Z_{N}\}$ and $Z_{i}$s are independent and identically distributed (i.i.d.) standard Gaussian distributed random variables. That is, $Z_{i}\sim\mathcal{N}(0,1)$. More specifically, $\Xi_{N}$ is the maximum of $N$ i.i.d. standard Gaussian random variables. In addition, $\Theta(N)$ is on the order of $\mathcal{O}(1)$, where $\mathcal{O}(\cdot)$ is the big $O$ notation. In particular, $\Theta(1)=0$. In this way, it holds that
$\mathbb{E}\{H_{\text{max}}\}=\frac{1}{v}\log\alpha+\frac{\sigma\sqrt{1-\frac{1}{\alpha^{2}}}}{\sqrt{2v^{3}}\delta}\left(\sqrt{\log N}+\Theta(N)\right),$ (23)
@@
Remark:  When $N=1$, it has been shown by Lemma 1 that the gradient noise does not affect the expectation of the first hitting time in the SGD algorithm. In particular, Eq. (23) also suggests the same conclusion. Further, when $N\geq 2$, the presence of the gradient noise introduces an additional term on the average first hitting time, which is proportional to the standard variance of the gradient noise and $\sqrt{\log N}$.
@@
Next, in order to make the problem analytically tractable and obtain insightful results, we consider the impact of gradient noise with the objective function $f(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{I}\boldsymbol{x}$. Thus, it holds that $\breve{\boldsymbol{x}}[j+1]=\breve{\boldsymbol{x}}[j]-\eta\breve{\boldsymbol{x}}[j]+\boldsymbol{w}.$ Note that $\eta$ is the learning rate and $\boldsymbol{w}$ is isotropic Gaussian noise, i.e., $\boldsymbol{w}\sim\mathcal{N}(0,\breve{\sigma}^{2}\boldsymbol{I})$. To simplify the notation, we define $\boldsymbol{r}[j]=\breve{\boldsymbol{x}}[j]-\boldsymbol{x}^{*}$ and $r=\left\|\boldsymbol{r}\right\|_{2}$. Due to the symmetry of the normal distribution, the projection of $\boldsymbol{w}$ onto the direction of $\boldsymbol{r}$, denoted by $\boldsymbol{w}_{\parallel}$, is also Gaussian distributed, i.e., $\left\|\boldsymbol{w}_{\parallel}\right\|_{2}\sim\mathcal{N}(0,\breve{\sigma}^{2})$. In this context, the component perpendicular to the $\boldsymbol{r}$ direction is given by $\boldsymbol{w}_{\perp}=\boldsymbol{w}-\boldsymbol{w}_{\parallel}$. In this case, we have $\langle\boldsymbol{w}_{\perp},\boldsymbol{r}\rangle=0$, in which $\langle\boldsymbol{a},\boldsymbol{b}\rangle$ denotes the inner product of the two vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. In addition, we find that $\mathbb{E}\{\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}\}=(N-1)\breve{\sigma}^{2}$. Thus, it holds that
$r[j+1]=r[j]-\eta r[j]+\left(\sqrt{r^{2}[j]+\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}-r[j]\right)+\boldsymbol{w}_{\parallel},$ (24)
@@
in which $w_{\parallel}=\left\|\boldsymbol{w}_{\parallel}\right\|_{2}$. For $r\gg\breve{\sigma}$, we have $\sqrt{r^{2}+\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}-r=\frac{\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}{2r}+\mathcal{O}(1/r^{2})$ by Taylor approximation $\sqrt{1+s}=1+\frac{1}{2}s+\mathcal{O}(s^{2})$ near $s=0$. For a given $r$, let the backward speed be defined as $z=\frac{\left\|\boldsymbol{w}_{\perp}\right\|_{2}^{2}}{2r}$ In this case, it can be seen that the presence of the noise $\boldsymbol{w}_{\perp}$ hinders the process of the gradient descent, the speed of which is given by $z$. More specifically, $z$ is Chi-square distributed with mean given by $\mathbb{E}\left\{z\right\}=\frac{(N-1)\breve{\sigma}^{2}}{2r}$, and variance given by $\text{Var}\left\{z\right\}=\frac{(N-1)\breve{\sigma}^{4}}{r^{2}}$. When $N=2$, the backward speed $z$ follows an exponential distribution. Therefore, the continuous-time approximation of Eq. (24) can be formulated as follows:
$\mathrm{d}r=-\left[\eta r-\frac{(N-1)\breve{\sigma}^{2}}{2r}\right]\mathrm{d}t+\sqrt{\sigma^{2}+\frac{(N-1)\breve{\sigma}^{4}}{r^{2}}}\mathrm{d}B(t).$ (25)
@@
The expectation of the first hitting time $H_{d}$ is given by $\mathbb{E}\{H_{d}\}=\mathcal{G}(r_{0})-{\mathcal{G}}(\delta)$, in which $\mathcal{G}(r)=\frac{1}{2\eta}\log\left[{\eta r^{2}-\frac{(N-1)\breve{\sigma}^{2}}{2}}\right]$.
@@
By integrating the primitive function ${\left[\eta r-\frac{(N-1)\breve{\sigma}^{2}}{2r}\right]^{-1}}$ with respect to $r$, and applying Lemma 1, we arrive at the corollary. ∎
In this way, one can notice that the presence of the gradient noise does not affect the expectation of the first hitting time of the gradient descent algorithm when $N=1$. Further, the above analysis also indicates that as the value of $r$ decreases, the adverse impact of gradient noise on convergence becomes increasingly severe. This also implies that as the training progresses, increasing the number of workers involved in training or enlarging the mini-batch size, both of which help reduce gradient noise, could accelerate the convergence of SGD. Note that a high-resolution quantization of the gradients with large gradient noise can lead to a lot of network bandwidth usage, thereby being inefficient due to the waste of communication resources.
@@ III-C The Convergence Speed Given Gradient Staleness
In this subsection, we present an analysis of SGD with stale gradients through an SDDE-based approach. For the sake of discussion, we first examine the case of fixed staleness, before extending our discussion to the case in which $\tau$ is a random variable. Here, let the gradient staleness in SDDE be denoted by a positive constant $\tau$. To understand the behavior of the solution to a deterministic differential equation, we substitute the sampling function $\varrho(t)=Ae^{{\lambda}t}$, in which ${\lambda}$ is a complex number, into the delay differential equation (15), we have the following characteristics equation:
@@
Hence, Eq. (26) has a single root when ${v}<0$. In this case, the SGD algorithm with a stale gradient overdamps and does not converge to the global minimum. In the following discussion, we shall first focus on the case $v>0$ before extending our discussion to the case $v\geq 0$. In particular, the characteristic equation has two negative real roots when $0<{v}\tau<\frac{1}{e}$. In this case, the SGD algorithm with a stale gradient decays monotonically to the global minimum of the quadratic objective function. Moreover, the characteristics equation has a pair of complex conjugate roots with negative real parts when $\frac{1}{e}<{v}\tau<\frac{\pi}{2}$. In this case, the SGD algorithm with stale gradient undergoes damped oscillations. When ${v}\tau>\frac{\pi}{2}$, the real parts of the roots of Eq. (26) becomes positive. Therefore, the SGD algorithm undergoes diverging oscillations. Note that $\eta$ is the step size of the SGD algorithm. According to the above discussion, we have the following theorem for the discrete-time SGD.
@@
When $v>0$, the two roots of the characteristics function Eq. (26) are given by
@@
Given the learning rate $\eta$ and the step staleness $\breve{\tau}$, the characteristic roots of the continuous-time approximation of the ASGD algorithm, $\breve{x}[j+1]=\breve{x}[j]-\eta v\breve{x}\left[{j-\breve{\tau}}\right]$, are given by $\lambda_{k}=\frac{W_{k}{(-{v}\eta\breve{\tau})}}{\eta\breve{\tau}}$ for $k=0,1$.
@@
With a fixed step size, $\lambda_{k}$ are real and negative in which $|\lambda_{0}|\leq|\lambda_{1}|$ when $\breve{\tau}<\frac{1}{e\eta{v}}$. In addition, it can be noticed that $|\lambda_{0}|$ increases as $\breve{\tau}$ increases for
$\breve{\tau}\in\left(0,\frac{1}{e\eta{v}}\right).$ (28)
This indicates that the presence of a small degree of gradient staleness accelerates the stochastic gradient descent algorithm, which aligns with the analysis in [16]. However, the SGD algorithm starts to oscillate when $\breve{\tau}\in(\frac{1}{e\eta{v}},\frac{\pi}{2\eta{v}})$. In addition, the real part of the roots becomes positive when $\breve{\tau}>\frac{\pi}{2\eta{v}}$. It indicates that, the SGD algorithm diverges with a large degree of gradient staleness.
Remark:  To reveal the characteristic roots of (15) in case $\tau$ is a random variable, we will present the characteristic roots with uniformly distributed gradient staleness in the forthcoming discussion. Consider a Gaussian distributed gradient staleness, i.e., $\tau\sim\mathcal{N}(\vartheta,\varsigma^{2})$. Numerically, we find that it holds that $\lambda_{k}\approx W_{k}(-v\vartheta)/\vartheta$ for $k=0,1$ when $\varsigma\ll\vartheta$ [^7] More specifically, we have the following observation of Gaussian distributed gradient staleness through numerical results. When $\tau\sim\mathcal{N}(\vartheta,\varsigma^{2})$ and $\varsigma\ll\vartheta$, ASGD nearly has the same convergence rate as when the gradient staleness is a constant, i.e., $\tau=\vartheta$. With a sufficiently large $\varsigma$, ASGD may diverge despite a small $\vartheta$. Analytical results for the staleness of Gaussian distribution gradients are noted as future work.